# Unleashing the Power of LLMs as Multi-Modal Encoders for Text and Graph-Structured Data

### v1

@@ 3.2. Fusion of Graph and Textual Data in LLM
3.3. Graph-Text Alignment via Contrastive Learning
Figure 2. Graph-text alignment through contrastive learning. Originally, each data input consists of a graph and its corresponding text. To align the graph space with the text space, each graph is converted into a language-based text description, leading to two processing streams: the original graph + text, and the new graph description + text. We then leverage contrastive learning where transformed versions of the same sample serve as positive pairs, while different samples within the batch act as negative pairs.
@@ 3.3. Graph-Text Alignment via Contrastive Learning
As we mentioned above, in Janus, graph embeddings are converted into the text space via a MLP module and then are processed by the LLM together with the input text tokens. To better align the graph embedding space and text space, we consider contrastive learning for the graph-text alignment.
@@
Contrastive learning has been broadly used in various domains (He et al., 2020; Chen et al., 2020; Lin et al., [n. d.], 2024; Yu et al., 2022; Iter et al., 2020a; Xu et al., 2022). The key idea behind contrastive learning is to minimize the distance between similar or positive pairs in the embedding space, while maximizing the separation between dissimilar or negative pairs. Although similar techniques have been applied in graph representation learning (Yu et al., 2022; You et al., 2020; Fang et al., 2024) and language models (Iter et al., 2020b; Fang et al., 2020; Giorgi et al., 2021), these approaches typically operate within a single modality (either graphs or text). In contrast, our Janus framework aligns the embedding spaces between text and graphs by (1) converting graph data into textual descriptions and (2) using contrastive learning to strengthen the connections between the textual descriptions of graphs and their corresponding original graph representations, as shown in Figure 2.
3.3.1. Translating Graphs into Textual Descriptions
@@
We first align the graph embeddings with text embeddings derived from descriptions of the same graph. Specifically, we start by listing all triples within the graph in a simple format: [entity (node), relation (edge), entity (node)]. This textual representation of the graph is then converted into a textual format by reorganizing the entities and relations into natural language. For example, the triples ”(analyzing, causes, new knowledge), (knowledge, causes, learn), (learn, causes, find information)” are transformed into the following textual description: ”analyzing causes new knowledge; knowledge causes learn; learn new causes find information.
3.3.2. Contrastive Learning for Alignment
@@
Now, we have two branches in the Janus framework: the original graph + text branch and the graph description + text branch. These branches create a dual-view architecture that facilitates the alignment of graph and text embedding spaces through contrastive learning.
@@
In the contrastive learning stage, we define positive pairs as the embeddings from the branch processing the original graph and its corresponding textual description that accurately reflect the same information. Negative pairs, on the other hand, consist of unrelated graph-text pairs from within the same batch, ensuring that the model learns to distinguish between semantically aligned and unaligned graph-text pair data. Specifically, for the original graph + text branch, we use the LLM backbone to generate the joint embeddings $\bm{z}_{\rm orig}$. Similarly, text embeddings $\bm{z}_{\rm new}$ for the graph description + text branch will be obtained from the LLM. By applying infoNCE loss (Oord et al., 2018), we have
@@
(6) $\small\mathcal{L}^{(\text{infoNCE})}=-\sum_{i=1}^{n_{\text{batch}}}\log\left(\frac{\exp\left(\tilde{\bm{z}}_{\rm orig}^{\mathrm{T}}(i)\cdot\tilde{\bm{z}}_{\rm new}(i)\right)}{\sum_{j=1}^{n_{\text{batch}}}\exp\left(\tilde{\bm{z}}_{\rm orig}^{\mathrm{T}}(i)\cdot\tilde{\bm{z}}_{\rm new}(j)\right)}\right)$
@@
where $n_{\text{batch }}$ represents the number of samples in a training batch, and $\tilde{\bm{z}}_{\text{orig}}$ and $\tilde{\bm{z}}_{\text{new}}$ denote the normalized vectors of $\bm{z}_{\text{orig}}$ and $\bm{z}_{\mathrm{new}}$, respectively. As we present in Section 4, the contrastive learning method improves Janus’s ability to better align the graph and text embeddings, which is further beneficial to the downstream tasks.
3.4. Training
Our proposed framework, Janus, can accommodate a wide array of tasks that require the integration of graph data and text. The training process for each task incorporates a task-specific loss function combined with a contrastive learning loss. The general form of the combined loss for each task can be represented as follows:
(7) $\mathcal{L}=\mathcal{L}^{(\rm task)}+\lambda\mathcal{L}^{(\text{infoNCE})}$
where $\lambda$ is a hyperparameter used to adjust the weights between loss functions. $\mathcal{L}^{(\rm task)}$ refers to the loss directly associated with the primary objective of the task. For KG-Contextualized QA, we use cross-entropy loss; for graph-text pair classification, we apply binary cross-entropy (BCE) loss; and for retrieval, we use infoNCE loss. More details can be found in the Appendix C.2.
4. Experiments
Next, we assess the performance of Janus and compare it to strong baselines on three types of tasks: KG-contextualized QA, graph-text pair classification, and retrieval tasks (§4.1-4.3). We also perform extensive ablation studies to understand the impact of our design and other choices (§4.4). We will make the code available upon acceptance.
4.1. KG-contextualized QA Performance
4.1.1. Datasets and Metrics
We first evaluated Janus on three QA datasets, i.e., CommonsenseQA (Talmor et al., 2019), OpenBookQA (Mihaylov et al., 2018), and MedQA-USMLE (Jin et al., 2020). We split these datasets according to (Yasunaga et al., 2021; Zhang et al., 2022) into train, validation, and test splits. For evaluation, we use accuracy as the metric to measure the performance on each dataset.
For each question in the QA datasets, a subgraph context extracted from a KG is utilized to provide additional contextual information, following Yasunaga et al. (Yasunaga et al., 2021). Specifically, for CommonsenseQA and OpeBookQA, we used the ConceptNet (Speer et al., 2017), which contains 799,273 nodes and 2,487,810 edges. Node embeddings are initialized by Roberta-Large (Liu et al., 2019) and kept frozen during the training process, following Yasunaga et al. (2021). The query node embeddings mentioned in Section 3.1 are calculated by the LLM backbone. For MedQA-USMLE dataset, we used the UMLS knowledge graph used in Yasunaga et al. (Yasunaga et al., 2021), which contains 9,958 nodes and 44,561 edges. Node embeddings are initialized using SapBERT (Liu et al., 2021), following Yasunaga et al. (2021).
4.1.2. Baselines
To validate the effectiveness of our Janus framework, we compare Janus with a range of baseline models. These baselines include both language models (LM) and hybrid approaches that integrate language models with knowledge graphs (LM+KG).
Fine-tuned LMs. We consider the following vanilla fine-tuned language models (LMs): RoBERTa-Large (Liu et al., 2019), E5-Mistral (Wang et al., 2023), and AristoRoBERTa (Clark et al., 2020). Additionally, for MedQA-USMLE dataset, we use several domain-specific models, including SapBERT (Liu et al., 2021) and BioBERT (Lee et al., 2020).
Existing LM+KG models. Our LM+KG baselines include: GreaseLM (Zhang et al., 2022) QAGNN (Yasunaga et al., 2021), Relation Network (RN) (Santoro et al., 2017), RGCN (Schlichtkrull et al., 2018), GconAttn (Wang et al., 2019), KagNet (Lin et al., 2019a), and MHGRN (Feng et al., 2020). Unlike these models that leverage GNNs as the backbone, our framework Janus adopts more powerful LLMs for better performance. We adopt E5-Mistral (Wang et al., 2023) as Janus’s LLM backbone.
4.1.3. Training Details
During the training process, we freeze the LLM part in Janus, only training GNNs and the MLP adapter. In this way, we reduced the computational resources required while still leveraging LLMs’ robust language understanding capabilities. For further details on the training process such as hyperparameters, please refer to the Appendix D.
4.1.4. Results
@@
Table 1. Test accuracy comparison on CommonsenseQA. The baseline results are mainly sourced from Zhang et al. (2022).
@@
RoBERTa-Large (Liu et al., 2019) 68.69 $\pm$ 0.56
E5-Mistral (Wang et al., 2023) 69.49 $\pm$ 0.28
@@
RGCN (Schlichtkrull et al., 2018) 68.41 $\pm$ 0.66
GconAttn (Wang et al., 2019) 68.59 $\pm$ 0.39
KagNet (Lin et al., 2019a) 69.01 $\pm$ 0.22
RN (Santoro et al., 2017) 69.08 $\pm$ 0.91
MHGRN (Feng et al., 2020) 71.11 $\pm$ 0.10
QA-GNN (Yasunaga et al., 2021) 73.41 $\pm$ 0.92
GreaseLM (Zhang et al., 2022) 74.20 $\pm$ 0.40
Janus (Ours) 81.09 $\pm$ 0.73
Table 2. Test accuracy comparison on OpenBookQA. The baseline results are mainly sourced from Zhang et al. (2022)
Methods Test Acc
Language Models Only
RoBERTa-Large (Liu et al., 2019) 64.80 $\pm$ 2.37
AristoRoBERTa (Clark et al., 2020) 78.40 $\pm$ 1.64
E5-Mistral (Wang et al., 2023) 74.80 $\pm$ 0.35
LM + KG + Scientific Facts
RGCN (Schlichtkrull et al., 2018) 74.60 $\pm$ 2.53
GconAttn (Wang et al., 2019) 71.80 $\pm$ 1.21
RN (Santoro et al., 2017) 75.35 $\pm$ 1.39
MHGRN (Feng et al., 2020) 81.87 $\pm$ 1.86
QA-GNN (Yasunaga et al., 2021) 82.77 $\pm$ 1.56
GreaseLM (Zhang et al., 2022) 83.87 $\pm$ 1.29
Janus (Ours) 86.67 $\pm$ 1.10
Janus (Ours) + Scientific Facts 93.33 $\pm$ 0.42
Table 3. Test accuracy comparison on MedQA-USMLE. The baseline results are mainly sourced from Zhang et al. (2022)
Methods Test Acc
Language Models Only
BERT-Base (Devlin et al., 2019) 34.3
BioBERT-Base (Lee et al., 2020) 34.1
RoBERTa-Large (Liu et al., 2019) 35
BioBERT-Large (Lee et al., 2020) 36.7
SapBERT (Liu et al., 2021) 37.2
E5-Mistral (Wang et al., 2023) 38.3
LM + KG
QA-GNN (Yasunaga et al., 2021) 38
GreaseLM (Zhang et al., 2022) 38.5
Janus (Ours) 49.9 $\pm$ 0.88
@@
We first conduct comparison experiments on CommonsenseQA and OpenBookQA datasets, as illustrated in Table 1 and 2. In both datasets, our framework outperforms all other methods significantly. Specifically, on CommonsenseQA, it surpasses the strongest baseline, GreaseLM, by a margin of 6.89%, achieving a test accuracy of 81.09%. This remarkable performance can be attributed primarily to the robust capabilities of the LLM backbone integrated within our framework. The LLM backbone not only enhances language understanding but also integrates the contextual knowledge derived from graphs, thereby substantially improving KG-contextualized QA performance. On OpenBookQA, Janus also demonstrates superior performance, achieving 86.67% test accuracy. When further integrating the extra corpus of scientific facts provided by OpenbookQA (Clark et al., 2020), which are leveraged by all the LM+KG baselines, our model reaches a remarkable 93.33% accuracy, exceeding that of all previous models that also utilized scientific facts by nearly 10%. These results clearly demonstrate the exceptional capability of Janus, with the LLM backbone playing a critical role in significantly enhancing both language understanding and the integration of external knowledge.
@@
We further evaluate Janus on the MedQA-USMLE dataset to assess its generalization capability across specialized domains and report the results in Table 3. Janus achieves a test accuracy of 49.9%, a substantial improvement over all the baselines, including domain-specific models. The significant boost in performance can be attributed to the LLM backbone’s ability to handle complex medical terminology and integrate knowledge from medical knowledge graphs effectively. This result further reinforces the adaptability of Janus across different domains showcasing its ability to excel not only in general and scientific question answering but also in knowledge-intensive and highly specialized fields.
In addition to evaluating Janus against various baseline models, we also conduct a direct comparison between Janus and its LLM backbone, E5-Mistral (Wang et al., 2023). For a fair comparison, we keep the E5-Mistral frozen, training only the classification head, which is consistent with how we evaluate Janus. This ensures that any performance differences can be attributed to the integration of graph knowledge rather than differences in backbone training. The experimental results demonstrate that Janus significantly outperforms the standalone E5-Mistral model across all datasets by 11-12%. These results clearly highlight the benefits of integrating structured knowledge graphs with the LLM backbone. While the LLM backbone alone captures rich linguistic information, the additional graph context enables the model to better understand the relationships and structured information relevant to the task, resulting in significant performance improvements across a wide range of QA tasks.
4.2. Graph-Text Pair Classification Performance
@@
Building upon the strong results achieved in the KG-contextualized QA tasks, we further evaluate Janus in a different task: graph-text pair classification. This task requires the model to determine whether a given text accurately corresponds to graph-structured data. Our objective in evaluating this task is to assess Janus’s ability to generalize beyond QA scenarios, demonstrating its versatility in handling a broader range of graph-text embedding tasks.
4.2.1. Datasets and Metrics
We use a dataset derived from WebNLG (Gardent et al., 2017), which contains 10,000 training, 4,000 validation, and 2,000 test graph-text pairs. Each graph consists of triples (entity, relation, entity) from DBPedia (Lehmann et al., 2015), and each pair is labeled as 1 if the text correctly describes the entities and their relationships, and 0 otherwise. Each of the train, validation, and test splits has an equal number of positive and negative samples. We use accuracy as the evaluation metric, measuring the proportion of correctly classified graph-text pairs. The details of how we curated this dataset can be found in the Appendix B.2. Unless otherwise specified, we refer to this dataset simply as WebNLG throughout the paper.
4.2.2. Baselines
We compare Janus against strong KG+LM baselines, including RGCN (Schlichtkrull et al., 2018), MHGRN (Feng et al., 2020), GreaseLM (Zhang et al., 2022), and QA-GNN (Yasunaga et al., 2021). These models have demonstrated strong performance in QA tasks and provide open-source implementations, allowing us to reliably compare their performance against Janus.
4.2.3. Results
Table 4. Test accuracy comparison on WebNLG dataset.
@@ 4.2.3. Results
RGCN (Schlichtkrull et al., 2018) 63.20 $\pm$ 0.49
MHGRN (Feng et al., 2020) 84.98 $\pm$ 0.53
QA-GNN (Yasunaga et al., 2021) 75.55 $\pm$ 3.54
GreaseLM (Zhang et al., 2022) 82.50 $\pm$ 4.29
Janus (Ours) 88.43 $\pm$ 1.33
The results of the graph-text pair classification task are presented in Table 4. Janus significantly outperforms all baseline models by 3.45%, achieving a test accuracy of 88.43%. While models like GreaseLM and MHGRN perform well by incorporating knowledge graphs, they lack the deeper contextual understanding due to the limitations of their less powerful backbone models, such as shallow transformers or GNNs.
4.3. Retrieval Performance
Next, we further evaluate Janus’s performance on retrieval tasks, where the objective is to retrieve relevant sentences or documents based on a query with graph-context.
4.3.1. Datasets and Metrics
We conduct experiments on two datasets, SciFact (Wadden et al., 2020) and FiQA (Maia et al., 2018), and report NDCG@10 as the primary evaluation metric.
4.3.2. Baselines
We compare Janus against a variety of baselines, including both language models alone (e.g., RoBERTa (Liu et al., 2019), E5 (Wang et al., 2022, 2023), GTR-XXL (Ni et al., 2022b)) and models that integrate knowledge graphs (e.g., QA-GNN (Yasunaga et al., 2021), GreaseLM (Zhang et al., 2022)).
4.3.3. Results
Table 5. Comparision results (NDCG@10) on retrieval tasks.

### v2

@@ 3.3 Graph-Text Alignment via Contrastive Learning
Figure 2: Overview of graph-text alignment through contrastive learning.
@@
As we mentioned above, in GT2Vec, graph embeddings are converted into the text space via a MLP module and then are processed by the LLM together with the input text tokens. To better align the graph embedding space and text space, we consider contrastive learning for the graph-text alignment.
@@
Contrastive learning has been broadly used in various domains He et al. (2020); Chen et al. (2020); Lin et al. ; Lin et al. (2024); Yu et al. (2022); Iter et al. (2020a); Xu et al. (2022). The key idea behind contrastive learning is to minimize the distance between similar or positive pairs in the embedding space, while maximizing the separation between dissimilar or negative pairs. Although similar techniques have been applied in graph representation learning Yu et al. (2022); You et al. (2020); Fang et al. (2024) and language models Iter et al. (2020b); Fang et al. (2020); Giorgi et al. (2021), these approaches typically operate within a single modality (either graphs or text). In contrast, our GT2Vec framework aligns the embedding spaces between text and graphs by (1) converting graph data into textual descriptions and (2) using contrastive learning to strengthen the connections between the textual descriptions of graphs and their corresponding original graph representations, as shown in Figure 2.
3.3.1 Translating Graphs into Textual Descriptions
@@
We first align the graph embeddings with text embeddings derived from descriptions of the same graph. Specifically, we start by listing all triples within the graph in a simple format: [entity (node), relation (edge), entity (node)]. This textual representation of the graph is then converted into a textual format by reorganizing the entities and relations into natural language. For example, the triples "(analyzing, causes, new knowledge), (knowledge, causes, learn), (learn, causes, find information)" are transformed into the following textual description: "analyzing causes new knowledge; knowledge causes learn; learn new causes find information."
3.3.2 Contrastive Learning for Alignment
@@
Now, we have two branches in the GT2Vec framework: the original graph + text branch and the graph description + text branch. These branches create a dual-view architecture that facilitates the alignment of graph and text embedding spaces through contrastive learning.
@@
In the contrastive learning stage, we define positive pairs as the embeddings from the branch processing the original graph and its corresponding textual description that accurately reflect the same information. Negative pairs, on the other hand, consist of unrelated graph-text pairs from within the same batch, ensuring that the model learns to distinguish between semantically aligned and unaligned graph-text pair data. Specifically, for the original graph + text branch, we use the way mentioned in §3.2 to generate the joint embeddings $\bm{z}_{\rm orig}$. This involves combining graph embeddings, obtained from a graph encoder, with textual embeddings generated by the LLM’s embedding layer to produce joint embeddings. For the graph description + text branch, we treat the input as pure text and use only the LLM to process it and generate the joint embeddings $\bm{z}_{\rm new}$. By applying infoNCE loss Oord et al. (2018), we have
@@
$\mathcal{L}^{(\text{infoNCE})}=-\sum_{i=1}^{n_{\text{batch}}}\log\left(\frac{\exp\left(\tilde{\bm{z}}_{\rm orig}^{\mathrm{T}}(i)\cdot\tilde{\bm{z}}_{\rm new}(i)\right)}{\sum_{j=1}^{n_{\text{batch}}}\exp\left(\tilde{\bm{z}}_{\rm orig}^{\mathrm{T}}(i)\cdot\tilde{\bm{z}}_{\rm new}(j)\right)}\right)$
@@
where $n_{\text{batch }}$ represents the number of samples in a training batch, and $\tilde{\bm{z}}_{\text{orig}}$ and $\tilde{\bm{z}}_{\text{new}}$ denote the normalized vectors of $\bm{z}_{\text{orig}}$ and $\bm{z}_{\mathrm{new}}$, respectively. As we present in §4, the contrastive learning method improves GT2Vec’s ability to better align the graph and text embeddings, which is further beneficial to the downstream tasks.
@@
Table 1: Test accuracy comparison on CommonsenseQA and OpenBookQA. The baseline results are mainly sourced from Zhang et al. (2022) and Yasunaga et al. (2021). Bold indicates the best result, and underline indicates the second best. LP means linear probe.
Methods CommonsenseQA OpenBookQA (w/o Scientific Facts) OpenBookQA (w/ Scientific Facts)
Language Models Only
RoBERTa-Large Liu et al. (2019) 68.69 $\pm$ 0.56 64.80 $\pm$ 2.37 78.40 $\pm$ 1.64
E5-Mistral, LP Wang et al. (2023) 69.49 $\pm$ 0.28 74.80 $\pm$ 0.35 81.67 $\pm$ 0.31
E5-Mistral, LoRA Wang et al. (2023) 78.73 $\pm$ 0.16 85.60 $\pm$ 0.20 91.87 $\pm$ 0.12
LM + KG
RGCN Schlichtkrull et al. (2018) 68.41 $\pm$ 0.66 62.45 $\pm$ 1.57 74.60 $\pm$ 2.53
GconAttn Wang et al. (2019) 68.59 $\pm$ 0.39 64.75 $\pm$ 1.48 71.80 $\pm$ 1.21
MHGRN Feng et al. (2020) 71.11 $\pm$ 0.10 66.85 $\pm$ 1.19 81.87 $\pm$ 1.86
QA-GNN Yasunaga et al. (2021) 73.41 $\pm$ 0.92 67.80 $\pm$ 2.75 82.77 $\pm$ 1.56
GreaseLM Zhang et al. (2022) 74.20 $\pm$ 0.40 65.60 $\pm$ 0.40 83.87 $\pm$ 1.29
GT2Vec, LP (Ours) 81.09 $\pm$ 0.73 86.67 $\pm$ 1.10 93.33 $\pm$ 0.42
GT2Vec, LoRA (Ours) 81.39 $\pm$ 0.11 88.13 $\pm$ 0.42 93.67 $\pm$ 0.31
3.4 Training
Our proposed framework, GT2Vec, can accommodate a wide array of tasks that require the integration of graph data and text. The training process for each task incorporates a task-specific loss function combined with a contrastive learning loss. The general form of the combined loss for each task can be represented as follows:
$\mathcal{L}=\mathcal{L}^{(\rm task)}+\lambda\mathcal{L}^{(\text{infoNCE})}$ (6)
where $\lambda$ is a hyperparameter used to adjust the weights between loss functions. $\mathcal{L}^{(\rm task)}$ refers to the loss directly associated with the primary objective of the task. For KG-Contextualized QA, we use cross-entropy loss; for graph-text pair classification, we apply binary cross-entropy (BCE) loss; and for retrieval, we use infoNCE loss. More details can be found in the Appendix D.2.
4 Experiments
Next, we assess the performance of GT2Vec and compare it to strong baselines on three types of tasks: KG-contextualized QA, graph-text pair classification, and retrieval tasks (§4.1-4.3). We also perform extensive ablation studies to understand the impact of our design and other choices (§4.4).
4.1 KG-contextualized QA Performance
4.1.1 Datasets and Metrics
We first evaluated GT2Vec on three QA datasets, i.e., CommonsenseQA Talmor et al. (2019), OpenBookQA Mihaylov et al. (2018), and MedQA-USMLE Jin et al. (2020). We split these datasets according to Yasunaga et al. (2021) and Zhang et al. (2022) into train, validation, and test splits. For evaluation, we use accuracy as the metric to measure the performance on each dataset. For each question in the QA datasets, a subgraph context extracted from a KG is utilized to provide additional contextual information, following Yasunaga et al. Yasunaga et al. (2021) (see Appendix C.1).
4.1.2 Baselines
We compare GT2Vec with a range of baseline models, including both language models (LM) and hybrid approaches that integrate language models with knowledge graphs (LM+KG).
Fine-tuned LMs. We consider the following vanilla fine-tuned language models (LMs): RoBERTa-Large Liu et al. (2019) and E5-Mistral Wang et al. (2023). Additionally, for MedQA-USMLE dataset, we use several domain-specific models, including SapBERT Liu et al. (2021) and BioBERT Lee et al. (2020).
Existing LM+KG models. Our LM+KG baselines include: GreaseLM Zhang et al. (2022) QAGNN Yasunaga et al. (2021), Relation Network (RN) Santoro et al. (2017), RGCN Schlichtkrull et al. (2018), GconAttn Wang et al. (2019), and MHGRN Feng et al. (2020). Unlike these models that leverage GNNs as the backbone, our framework GT2Vec adopts more powerful LLMs for better performance. We adopt E5-Mistral Wang et al. (2023) as GT2Vec’s LLM backbone.
4.1.3 Training Details
During training, we adopt parameter efficient finetuning techniques, including Linear Probe (LP) Radford et al. (2021) and Low-Rank Adaptation (LoRA) Hu et al. (2022), for E5-Mistral and GT2Vec. LP freezes the pre-trained LLM and trains a linear classifier on top, while LoRA introduces low-rank adaptation layers for efficient finetuning. we employ full parameter fine-tuning for the other baselines. Further details on the training process such as hyperparameters can be found in Appendix E.
Table 2: Test accuracy comparison on MedQA-USMLE. The baseline results are mainly sourced from Zhang et al. (2022). Bold indicates the best result, and underline indicates the second best. LP means linear probe.
@@ 4.1.3 Training Details
BERT-Base Devlin et al. (2019) 34.3
BioBERT-Base Lee et al. (2020) 34.1
RoBERTa-Large Liu et al. (2019) 35
BioBERT-Large Lee et al. (2020) 36.7
SapBERT Liu et al. (2021) 37.2
E5-Mistral, LP Wang et al. (2023) 39.4 $\pm$ 1.1
E5-Mistral, LoRA Wang et al. (2023) 51.1 $\pm$ 0.3
@@
QA-GNN Yasunaga et al. (2021) 38
GreaseLM Zhang et al. (2022) 38.5
GT2Vec, LP (Ours) 49.9 $\pm$ 0.9
GT2Vec, LoRA (Ours) 53.4 $\pm$ 0.3
4.1.4 Results
@@
We first conduct comparison experiments on CommonsenseQA and OpenBookQA datasets, as illustrated in Table 1. In both datasets, our framework outperforms all other methods significantly. Specifically, on CommonsenseQA, it surpasses the strongest baseline E5-Mistral, achieving a test accuracy of 81.39%. On OpenBookQA, GT2Vec also demonstrates superior performance, achieving 88.13% test accuracy. When further integrating the extra corpus of scientific facts provided by OpenbookQA Clark et al. (2020), our model reaches a remarkable 93.67% accuracy, outperforming all the baseline models that also utilized scientific facts. The remarkable performance can be attributed primarily to the robust capabilities of the LLM backbone integrated within our framework. The LLM backbone not only enhances language understanding but also integrates the contextual knowledge derived from graphs, thereby substantially improving KG-contextualized QA performance.
@@
We further evaluate GT2Vec on the MedQA-USMLE dataset to assess its generalization capability across specialized domains and report the results in Table 2. GT2Vec achieves a test accuracy of 53.4%, outperforming all the baselines including domain-specific models. This result further reinforces the adaptability of GT2Vec across different domains showcasing its ability to excel not only in general and scientific question answering but also in knowledge-intensive and highly specialized fields.
4.2 Graph-Text Pair Classification Performance
@@
Building upon the strong results achieved in the KG-contextualized QA tasks, we further evaluate GT2Vec in a different task: graph-text pair classification. Our objective in evaluating this task is to assess GT2Vec’s ability to generalize beyond QA scenarios, demonstrating its versatility in handling a broader range of graph-text embedding tasks. We evaluate models on the WebNLG Gardent et al. (2017) dataset and use accuracy as the evaluation metric, measuring the proportion of correctly classified graph-text pairs.
Table 3: Test accuracy comparison on WebNLG dataset.
@@
RGCN Schlichtkrull et al. (2018) 63.20 $\pm$ 0.49
MHGRN Feng et al. (2020) 84.98 $\pm$ 0.53
QA-GNN Yasunaga et al. (2021) 75.55 $\pm$ 3.54
GreaseLM Zhang et al. (2022) 82.50 $\pm$ 4.29
GT2Vec, LP (Ours) 88.43 $\pm$ 1.33
GT2Vec, LoRA (Ours) 89.70 $\pm$ 0.34
The results of the graph-text pair classification task are presented in Table 3. GT2Vec significantly outperforms all baseline models by at least 4.72%, achieving a test accuracy of 89.70%. While models like GreaseLM and MHGRN perform well by incorporating knowledge graphs, they lack the deeper contextual understanding due to the limitations of their less powerful backbone models, such as shallow transformers or GNNs.
4.3 Retrieval Performance
Next, we further evaluate GT2Vec’s performance on retrieval tasks, where the objective is to retrieve relevant sentences or documents based on a query with graph-context. We conduct experiments on two datasets, SciFact Wadden et al. (2020) and FiQA Maia et al. (2018), and report NDCG@10 as the primary evaluation metric.
Table 4: Comparision results (NDCG@10) on retrieval tasks.